OpenStudy (anonymous):

If a circle is circumscribed around a square and we know the radius of the circle how do we find the are of the square?

OpenStudy (anonymous):

A = pi r^2

OpenStudy (anonymous):

I know how to find the area of the circle, but I need the area of the square inside

OpenStudy (amistre64):

the diagonal of the square is equal to the diameter of the circle

OpenStudy (amistre64):

use the pythag thrm to find the length of the sides

OpenStudy (amistre64):

s^2 + s^2 = diameter^2 2s^2 = d^2 d^2 s^2 = --- 2 d^2 s = sqrt( --- ) 2 s = d/sqrt(2)

OpenStudy (amistre64):

or simply stop at s^2 since the area of a square is s*s :)

OpenStudy (anonymous):

Thanks so much! Cheers

OpenStudy (anonymous):

Something to ponder...